i need to know how to get y^2-12y-x^2-2x= -31 in standard form

Answer by stanbon(75887)

y^2-12y-x^2-2x= -31 in standard form
-----
Complete the square on the x-terms and on the y-terms.
----
y^2 - 12y + (12/2)^2 - (x^2 + 2x + (2/2)^2) = -31 + (12/2)^2 - (2/2)^2
-----
Factor:
(y-6)^2 - (x+1)^2 = 4
------
[(y-6)^2 / 4] - [(x+1)^2/ 4] = 1
=======================
Cheers,
Stan H.
=======================